Find all the units of Z[ √-7].


1
Expert's answer
2021-04-28T16:32:05-0400

Let x,y "\in" Z[-7]

Then x has the form a+b"\sqrt{-7}" and y has the form c+d"\sqrt{-7}" for arbitrary x and y.

"If\\
(a+b\sqrt{-7})(c+d\sqrt{-7})=1 , then\\
(a-b\sqrt{-7})(c-d\sqrt{-7})=1, so
1=(a+b\sqrt{-7})(c+d\sqrt{-7})(a-b\sqrt{-7})(c-d\sqrt{-7})=1\\
\text{which gives}\\
1=(a^2+7b^2)(c^2+7d^2)\\
So,\\
b=d=0,and \\ 1=a^2c^2,\\
\text{so either}\\
a=c=1, or, a=c=-1\\
\text{and these are the only units in}Z\sqrt{-7}"